Question 19 (5 points) Solve 5 g -2) + g = 6(g - 4). a) all numbers b) 0 c)2 O d) 0
Answer #1

fol: Ginen equation 5(8-2) +9 =6(9-4) 59-10+9=69-24 69-10 69-24 - 10 =-24 is not possible) No value for ģ which satsches the